Fermi National Laboratory
 

Accelerator Update for June 22, 2006 - June 23, 2006

 

Thursday June 22

The day shift began with Operations monitoring a Recycler stash, with the Tevatron (TeV) recovering from a quench, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for colliding beam, with the Antiproton Source (Pbar) stacking, with the Main Injector (MI) and Recycler conducting studies, and with MiniBooNE and NuMI taking beam.

Operations put the TeV into a dry squeeze at 9:34 AM.

Operations began preparing the Recycler at 9:39 AM to stash antiprotons. The stash ended and Pbar resumed stacking at 10:17 AM.

Operations added an event at 10:21 AM to the time line for Switchyard (SY) experts to conduct a study. They eventually got beam to F17.

TeV experts reported at 2:10 PM that they understood the problem with the orbit stabilization program. Experts completed a kicker study and then began a chromaticity study.

The Recycler began preparing to stash at 6:32 PM. The stash ended and Pbar resumed stacking at 7:40 PM.

TeV experts completed their studies at 7:42 PM.

Operations put the TeV into shot setup at 8:38 PM.

Operations established store 4794 at 11:40 PM with an initial luminosity of 61.4E30.

Friday June 23

The midnight shift began with Operations monitoring store 4794 and the stash, with Pbar stacking, with normal MI, Recycler, SY, and Booster conducting studies, and with MiniBooNE and NuMI taking beam.

The Recycler began preparing to stash at 1:36 AM.

At 2:19 AM, the NuMI beam permit tripped off on a target LCW flow indication. Ten minutes later, a NuMI expert directed operators to reset the beam permit and run at a lower intensity.

The stash ended and Pbar resumed stacking at 2:37 AM.

The Recycler began preparing to stash at 7:10 AM. The stash ended and Pbar resumed stacking at 7:56 AM.

The Plans for Friday and the Weekend

The plans for today are to allow TeV experts to conduct end-of-store studies. Booster experts will conduct two two-hour studies: one destructive and the other at low intensity. The weekend will consist of stacking, stashing, and storing.

Machine Reports

Tevatron

Problems

  • Flying wires still report sigmas okay, but the data calculated by OAC that fills the reports are still invalid
  • D0 reported that proton halo losses are higher than 2005, found that D0 vacuum is worse

Work Done This Week

  • Have CDF and D0 IP position changes ready to implement waiting for an end of store
  • Measured alpha bumps at CDF and D0 and implemented a vertical change at D0 of ~ 20cm
  • Tested new application that can measure coupling and calibrated SQ and SQA0 mults used for shot setup tuning
  • Investigated why the C100, chromaticity measurement application used in shot setup is not working

Work For Next Week

  • Need to move positions for CDF and D0

Linac

Last Week Summary

  • Running on the H- Ion Source
    • Beam Current 30 to 35 mA
  • Beam positions Started Out Normal, but After Tank 5 Quad Shorted they Retuned, which Resulted in better positions
  • Downtime 5.94 Hours With 31 Occurrences
    • 94.1% Availability
  • Major Sources of Downtime
    • Klystron 4 Phase Problems .64 hr
    • Tank 5 Quad #20 Shorted to Ground 3.94 hr

Problems

  • TRIM90 current was wandering and we replaced the supply
  • Debuncher Water Skid Tripped Several Times

Booster

Summary

  • Continued to work on orbits
  • Added second notcher (L12 kicker)
  • Changed 400 MeV control processor
    • Trying to speed up power supply controls
  • Worked on dampers
    • Quad damper on, mode 49 still off
  • Gamma –t and sum resonance studies
  • Might start collimator work today

First Notcher

Second Notcher

Next Week

  • Secondary Collimators (starting today)
    • Scans
    • Fix positions
  • Dampers will get tuned with higher intensity
  • RF work – several weak stations
  • Beam position survey

Main Injector

MI

  • I68 problem.
    • 120 GeV antiproton production
    • 120 Slow spill

MI BPM upgrade

  • MI30 BPMs are converted to run on new system.

New BLM system

  • Still trying to debug the new system at MI40
  • The BLMs in MI30 were still with multibus system.

Antiproton Source

Last Week

  • First “routine” week since shutdown. Nearly entire week spent stacking, tuning
  • Major downtime ~ 1 shift (Mon owl) for M:HV102 supply
  • Best stacking rate ~ 13 E10/hr (vs. 20 in February)
  • Beam on target ~ 5.5E12 (vs. full intensity of 8.0)

Things To Do

  • Yield to end of AP2 line same as before shutdown
    • Need to streamline/calibrate intermediate stacking efficiencies for Debuncher, D→A, etc.
  • We need to measure A:PRDCTN vs. beam on target and vs. cycle time, to understand our bottlenecks
    • Get PRDCTN back to 21 and keep it there at full rate
  • Access task list growing: Deb tank, CV1, QS732, TOR724, VT726, VT716, M:Q209, AP1 OTR
  • Would be fantastic to have hard-wired Ethernet connection to F23 and F27 for BPM readout (AP1, AP2, AP3)

MiniBooNE

Horn Supply Failure

  • Began about 22:00, Sunday. Experts called in.
  • Traced to arcing in supply. This damaged the connection between two Stripline busses. Shut down.
  • Monday, EE support removed damaged components and fabricated replacements
  • Tuesday, repaired horn and did thermal imaging
  • Smooth running since then

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/23/2006    email Fermilab